40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 9 10:52:37 2014, manasa, Update, General, Power outage II & recovery 
    Reply  Thu Oct 9 11:56:35 2014, Steve, Update, VAC, Power outage II & recovery 
       Reply  Thu Oct 9 13:29:14 2014, Jenne, Update, PSL, Power outage II & recovery 
       Reply  Thu Oct 9 17:33:28 2014, Steve, Update, VAC, Power outage II & recovery pumpingAgain.png
          Reply  Thu Oct 9 19:14:04 2014, ericq, Update, General, Power outage II & recovery 
          Reply  Fri Oct 10 14:41:04 2014, Steve, Update, VAC, Power outage II & recovery recovery_poweroutage.png
Message ID: 10588     Entry time: Thu Oct 9 13:29:14 2014     In reply to: 10587
Author: Jenne 
Type: Update 
Category: PSL 
Subject: Power outage II & recovery 

Quote:

 

 IFO vacuum, air condition and PMC HV are still down. PSL out put beam is blocked on the table.

 PMC is fine.  There are sliders in the Phase Shifter screen (accessible from the PMC screen) that also needed touching. 

PSL shutter is still closed until Steve is happy with the vacuum system - I guess we don't want to let high power in, in case we come all the way up to atmosphere and particulates somehow get in and get fried on the mirrors. 

ELOG V3.1.3-